Each interior angle of a regular polygon is 160. Fin the number of its sides and sum of interior angle of the polygon
Answer by fcabanski(1391)

a + 360/n = 180. a is the size of the interior angles of the regular polygon, and n is the number of sides.


160 + 360/n = 180


360/n = 20


n = 360/20 = 18 sides.


The sum of interior angles is 18 * 160 = 2880 degrees.